Circular No. 8855

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 N1 (McNAUGHT) R. H. McNaught reports his discovery of a comet on CCD images

obtained with the 0.5-m Uppsala Schmidt telescope at Siding Spring (discovery observation tabulated below); stacked images show the object to be slightly diffuse with a diffuse tail ∼ 10′′ long in p.a. 240^{o}. Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, A. C. Gilmore writes that his CCD images taken on July 11.64–11.70 UT with the Mt. John 1.0-m f/7.7 reflector show a small, condensed coma and a short fan tail in p.a. 240^{o}. S. Casulli notes that his CCD images taken on July 12.066 with a 0.40-m f/4.5 reflector at the Osservatorio Astronomico Vallemare di Borbona show a coma nearly 12′′ in diameter, elongated toward p.a. 80^{o}.

2007 UT  α_{2000}  δ_{2000}  Mag.
July 10.73325  1^{h}42^{m}19.^{s}69  +1^{o}10′08′′.4  17.7

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-N29.

T = 2007 Sept. 7.236 TT  ω = 266.^{o}796
Ω = 115.220 }2000.0
q = 2.28009 AU  i = 9.329

COMET C/2006 VZ_{13} (LINEAR) M. L. Sitko, University of Cincinnati and Space Science Institute; L.

Beerman, University of Cincinnati; R. W. Russell, D. K. Lynch, and R. Pearson, The Aerospace Corporation; H. B. Hammel, Space Science Institute; and W. Golisch, Infrared Telescope Facility (IRTF), NASA, report on observations made of comet C/2006 VZ_{13} using the IRTF (+ BASS) on July 9 UT. The comet exhibited a continuum between 8 and 13 μm, on top of which a silicate emission band from 8.5 to 12.2 μm was observed. An underlying blackbody, normalized to the continuum fluxes at 8.1 and 12.5 μm, yielded a mean grain temperature of 275 K (estimated uncertainty ± 5 K). The derived temperature was 6 percent higher than that of an equivalent radiative equilibrium blackbody at the heliocentric distance of the comet. The feature-to-continuum ratio in the silicate band was 1.27. The measured flux between 10.0 and 11.0 μm, using the 3′′.4 circular entrance aperture of BASS, was 1.6 ± 0.2 Jy (equivalent magnitude N = 3.5 ± 0.1).

Visual m_{1} estimates by J. Gonzalez, Leon, Spain: June 12.05 UT, 10.3 (25×100 bin.); 23.03, 9.4; 26.11, 8.8; July 4.91, 7.9 (7×50 bin.); 10.01, 7.3.

2007 July 12  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green